# Principal stratification with $U$-statistics under principal ignorability

### v3

@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 1, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
5.1 The probability index estimands with an intermediate variable
[…]
$D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\},$
$Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1),$
for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. In this simulation study, we consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance), and assume the interest lies in estimating all three principal probability index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
@@ 5.1 The probability index estimands with an intermediate variable
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions). Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
@@
Figure 1: Violin plots of estimators for the principal probability index estimand in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
5.2 The win ratio estimands with an intermediate variable
@@
In the second simulation study, we consider the case of the win ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
$\log\frac{\mathbb{P}(Y_{i}(Z_{i})\leq q)}{\mathbb{P}(Y_{i}(Z_{i})>q)}=\eta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$
@@
for $q\in\{1,2\}$, where $\eta_{1,0}=-1$ and $\eta_{2,0}=1$. Transformed covariates are created following the procedure described in the previous simulation study. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated via the ordered logit regression, while treatment propensity and principal scores are similarly estimated via logistic regression. For the debiased machine learning estimator, we used the multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean, and the treatment propensity and principal scores are estimated using the SuperLearner package. The debiased machine learning estimator adopts a five-fold cross-fitting ($K=5$ and $L=15$) procedure. We set the number of observations $n=2,000$ with $1,000$ replications. Violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure 2 in Section S6 of the Online Supplement.
@@
From Figure 2, we observe that, similar to the estimation of the probability index estimand in Section 5.1, the plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (shown in Web Figure 2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the rest in terms of estimation accuracy (bias).
@@
Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment probability or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment probability (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). Horizontal dashed lines represent the true values of the PGCEs.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The dataset (named ‘JC’ with $9,240$ observations, available in the R package causalweight) was from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, i.e., receiving an offer to participate in Job Corps immediately, or, the control group, i.e., not participating in Job Corps until three years later. Follow-up surveys were administered to both groups 24, 36, and 48 months after the random assignment to collect data on the outcomes of interest. As in many open-label randomized controlled trials, substantial noncompliance to treatment assignment as an intermediate outcome occurred 1 year after randomization (15.4% [$857/5,577$] of participants assigned to treatment did not adhere to the assignment, and 50.6% [$1,854/3,663$] assigned to control switched to the treatment condition).
@@
The proposed methods were implemented to examine the causal effect of Job Corps intervention, where the weekly earnings one year after the assignment were selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including baseline characteristics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 The SPI for continuous outcomes
@@
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al. (2009). Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
@@
The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.
@@
Table 1: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (14); and ‘DML’ for the debiased machine learning estimators given in (15). The corresponding 95% confidence interval (CI) is given in the following parenthesis.
@@
6.2 Principal win ratio for ordinal outcomes

### v4

@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for the PGCEs, including the three plug-in moment estimators motivated by the nonparametric identification results in Section S2 of the Online Supplement, the triply robust estimators with parametric working models given in (13), and the debiased machine learning estimators given in (14), under different data-generating processes.
5.1 Probabilistic index with an intermediate variable
In the first simulation study, we consider the case of the probabilistic index estimand introduced in Examples 1 and 3, where the potential outcomes are continuous and drawn from a normal distribution. We first generate four pre-treatment covariates $X_{i}=(X_{i1},X_{i2},X_{i3},X_{i4})^{\prime}$, where $X_{i1},X_{i2},X_{i3}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$, and $X_{i4}\overset{\text{i.i.d.}}{\sim}\mathcal{B}(0.5)$. Then, we draw the treatment assignment from $Z_{i}|X_{i}\sim\mathcal{B}\{\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.1X_{i4})\}$. Finally, the potential intermediate variable and the potential outcome are drawn from $D_{i}(Z_{i})|Z_{i},X_{i}\sim\mathcal{B}\{\mathrm{expit}(-1+2Z_{i}+X_{i1}-0.8X_{i2}+0.6X_{i3}-X_{i4})\}$, and $Y_{i}(Z_{i})|D_{i},Z_{i},X_{i}\sim\mathcal{N}(10+2Z_{i}-D_{i}+8X_{i1}+6X_{i2}+9X_{i3}+7X_{i4},1)$ for $i\in\{1,\ldots,n\}$, where, $\mathrm{expit}(x)=1/\{1+\exp(-x)\}$. We consider a binary intermediate variable ($D$ can be considered as the treatment receipt in the context of non-compliance) and assume the interest lies in estimating all three principal probabilistic index estimands ($\tau_{h}^{10}$, $\tau_{h}^{00}$ and $\tau_{h}^{11}$ with $h(u,v)=\mathbbm{1}(u\geq v)$). In addition, we create transformed covariates, $\widetilde{X}_{i}=(\widetilde{X}_{i1},\widetilde{X}_{i2},\widetilde{X}_{i3},\widetilde{X}_{i4})^{\prime}$, where
@@ 5.1 Probabilistic index with an intermediate variable
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. It is worth noting that, although in theory, machine learning methods are flexible with richer model spaces as $n$ increases, model misspecification can still be relevant to the debiased machine learning estimators, as their estimation accuracy in finite samples can be affected by the quality of the input covariate feature matrix (Cheng and Li, 2025).
For the plug-in moment estimators and the triply robust estimator, treatment propensity scores and principal scores are estimated using logistic regression, whereas the pairwise outcome mean is estimated using linear regression, as in Example 3. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear model libraries. We set the number of observations to $n=2,000$ and consider $1,000$ replications. Violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure S1 in Section S6 of the Online Supplement.
Figure 1 summarizes the results of the five estimators in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios, regardless of whether all nuisance functions or subsets of nuisance functions are correctly specified. Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure S1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In contrast, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator yields a slight improvement, as it is generally less biased.
@@
Figure 1: Violin plots of estimators for the principal probabilistic index estimand in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14). Horizontal dashed lines represent the true values of the PGCEs.
5.2 Win ratio with an intermediate variable
@@
We further consider the case of estimating the win ratio estimands introduced in Examples 2 and 4, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,2,3\}$, we consider the following three-category ordinal logistic regression model:
$\log\frac{\mathbb{P}\{Y_{i}(Z_{i})\leq q\}}{\mathbb{P}\{Y_{i}(Z_{i})>q\}}=\zeta_{q,0}+2Z_{i}-D_{i}+X_{i1}-X_{i2}+1.2X_{i3}-0.8X_{i4},$ (16)
@@
for $q\in\{1,2\}$, where $\zeta_{1,0}=-1$ and $\zeta_{2,0}=1$. Transformed covariates are created following Section 5.1. For the plug-in moment estimators and the triply robust estimator with parametric working models, the pairwise outcome mean is estimated using ordered logit regression, while treatment propensity and principal scores are estimated using logistic regression. For the debiased machine learning estimator, we used multinomial generalized linear models with the elastic net penalty from the glmnet package in R to estimate the pairwise outcome mean function. The treatment propensity and principal scores were estimated using the SuperLearner. The debiased machine learning estimator employs a five-fold cross-fitting procedure with $K=5$ and $L=15$. We set the number of observations $n=2,000$ with $1,000$ replications. In parallel to Section 5.1, violin plots summarizing the simulation results of each considered estimator are given in Figure 2 and Web Figure S2 in Section S6 of the Online Supplement.
@@
From Figure 2, we observe that the plug-in moment estimators are consistent whenever all relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent in all four scenarios where at most one of the three nuisance functions is misspecified. When at least two of the three nuisance functions are misspecified (as shown in Web Figure S2 in Section S6 of the Online Supplement), all considered estimators are biased, with the triply robust and debiased machine learning estimators outperforming the others in terms of estimation accuracy (bias).
@@
Figure 2: Violin plots of estimators for the win ratio estimands in scenarios where at most one nuisance function is misspecified. Labels: ‘tp’ for treatment propensity or propensity score; ‘ps’ for principal score; ‘oc’ for pairwise outcome mean; ‘tp+ps+oc’ for the scenario where all three nuisance functions are correctly specified, and ‘tp+ps’ for the scenario where the treatment propensity (propensity score) and the principal score are correctly specified, etc; ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14). Horizontal dashed lines represent the true values of the PGCEs.
5.3 Additional simulations under non-monotonicity
We consider a third simulation study to demonstrate the potential bias in the uncorrected triply robust and debiased machine learning estimators, as well as to validate the bias-corrected estimators under non-monotonicity. To achieve these goals, we modify the data-generating processes in Sections 5.1 and 5.2 by integrating the sensitivity function $\eta(X)$. Specifically, continuing with $p_{z}(X)=\mathrm{expit}(-1+2z+X_{1}-0.8X_{2}+0.6X_{3}-X_{4})$ for $z\in\{0,1\}$, we adopt the following configuration for the sensitivity function: $\eta(X)=\eta_{0}[1-\{p_{1}(X)-p_{0}(X)\}/\min\{p_{1}(X),1-p_{0}(X)\}]$, where, we let $\eta_{0}\in[0,1]$ to ensure the true sensitivity function is well-defined within the permissible bound. We consider two scenarios: one with moderate violations, represented by $\eta_{0}=0.5$, and another with severe violations, represented by $\eta_{0}=0.99$. The strata membership $S$ is drawn from from a multinominal distribution $\mathcal{M}\mathcal{N}\{e_{10,\eta}(X),e_{01,\eta}(X),e_{00,\eta}(X),e_{11,\eta}(X)\}$, where $\{e_{s,\eta}(X),s\in\{10,01,00,11\}\}$ are obtained following (15). The treatment assignment $Z\sim\mathcal{B}\{\mathrm{expit}(-X_{1}+0.5X_{2}-0.25X_{3}-0.1X_{4})\}$, and the intermediate variable $D$ is determined once $Z$ and $S$ are known. That is, $D=1$ if $Z=1,S\in\{10,11\}$ or $Z=0,S\in\{01,11\}$, and $D=0$ if $Z=1,S\in\{01,00\}$ or $Z=0,S\in\{10,00\}$. The potential outcome $Y(z)$ for $z\in\{0,1\}$ is generated following (16), based on the same covariates simulated as in Section 5.1.
To focus on the main ideas in this sensitivity simulation, we simplify without further considering nuisance function misspecification. We compare four estimators: the triply robust and debiased machine learning estimators under monotonicity, and the bias-corrected estimators with a known and correctly specified $\eta(X)$. The results are presented in Web Figure S3 in Section S6 of the Online Supplement.
Overall, the simulation results show that the bias-corrected estimators give consistent estimates of all four PGCEs under both non-monotonicity scenarios. As $\eta_{0}$ increases from $0.5$ to $0.99$, the biases of the triply robust and debiased machine learning estimators increase accordingly, due to a higher degree of non-monotonicity. Interestingly, such increases in bias are not uniform across all PGCEs. For example, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{10}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{10}$ are relatively small under both $\eta_{0}$ values, suggesting that they may be dramatically affected by violations of the monotonicity assumption in our setup. On the other hand, the biases of $\widehat{\tau}_{h,\mathrm{tr}}^{00}$, $\widehat{\tau}_{h,\mathrm{dml}}^{00}$, $\widehat{\tau}_{h,\mathrm{tr}}^{11}$, and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ increased significantly, indicating their relatively higher sensitivity to non-monotonicity. In further exploration, we find that as $\eta_{0}$ increases, a larger proportion of the $01$ stratum comes from the $00$ and $11$ strata, hence contributing to increased bias in the estimators for $\tau_{h}^{00}$ and $\tau_{h}^{11}$.
@@ 6 Data example
We illustrate the proposed methods with a data example from the U.S. Jobs Corps study (Schochet et al., 2001), where individualized academic education, vocational training, counseling, and job placement assistance were offered at Jobs Corps centers to economically disadvantaged youth aged from 16 to 24. The data were from a randomized controlled trial conducted at 119 Job Corps centers across the U.S., where eligible youth were randomly assigned to either the treatment group, which received an immediate offer to participate in Job Corps, or the control group, which did not participate in Job Corps until three years later. As in many open-label randomized controlled trials, substantial noncompliance with treatment assignment occurred as an intermediate outcome 1 year after randomization: 15.4% ($857/5,577$) of participants assigned to treatment did not adhere to the assignment, and 50.6% ($1,854/3,663$) assigned to control switched to the treatment condition.
@@
The proposed methods were implemented to examine the causal effect of the Job Corps intervention, with weekly earnings one year after the assignment selected as the final outcome. As the earning outcomes are often skewed and non-normal, studying alternative estimands that are not based on the mean contrasts may provide complementary insights into the treatment effects. In each data analysis below, we adjust for baseline covariates, including demographics (e.g., age, gender, ethnicity, etc.), education level, receipt of public assistance, involvement with the criminal justice system, crimes committed against Job Corps participants, use of legal and illegal substances, health and mortality, family formation and child care, and perform principal stratification analysis under the assumption that mean principal ignorability holds conditional on these covariates.
6.1 SPI for continuous outcomes
@@
We first implement the proposed methods by considering the job status one year after the assignment as the intermediate outcome, $D$, where if participant $i$ is employed one year after the assignment, then $D_{i}=1$ and the final outcome $Y_{i}$, the weekly earning, can be observed and defined without ambiguity; otherwise, $D_{i}=0$ and $Y_{i}$ is not well-defined and denoted by $Y_{i}=*$. In other words, we focus on the initial random assignment (an intention-to-treat analysis) and address earnings outcomes truncation by “death” (employment status); also see Zhang et al. (2009) for a similar setup. Under this setting, we estimate the SPI estimand defined in (3) for the always-employed (principal stratum $11$), i.e., participants who will be employed with or without Jobs Corps. For methods based on parametric working models, a logistic model is considered for the employment process within each assignment group, and a log-normal model is fitted for the final earning outcome to induce the pairwise mean outcome function. For the debiased machine learning approach, we use a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner with the random forest and generalized linear models libraries.
@@
The estimated results for SPI are presented in Table 2, where $1,000$ bootstrap replicates were generated to estimate the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probabilistic index is estimated to be $0.539$ using the triply robust and debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates obtained under the plug-in moment methods are slightly closer to the null compared to those from the triply robust and debiased machine learning methods, although their 95% confidence intervals continue to exclude the null.
@@
Table 2: Estimated SPI for the Jobs Corps data. ‘M1’, ‘M2’, and ‘M3’ for plug-in moment estimators $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, respectively; ‘TR’ for the triply robust estimators given in (13); and ‘DML’ for the debiased machine learning estimators given in (14). The corresponding 95% confidence interval (CI) is given in the following parentheses.
@@
6.2 Principal win ratio (PWR) for ordinal outcomes